Determine whether each system of linear equations has a) one and only one solution, b) infinitely many solutions, or c) no solution. Find all solutions whenever they exist.
2x-4y=5
3x+2y=6

Answer by mananth(16946)

2 x -4 y = 5 .............1
3 x + 2 y = 6 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 2
2 x -4 y = 5
6 x + 4 y = 12
Add the two equations
8 x = 17
/ 8
x = 2.13
plug value of x in (1)
2 x + -4 y = 5
4.25 + -4 y = 5
-4 y = 5 -4.25
-4 y = 0.75
y = -0.19
They have only one solution ( 2.13,-0.19)